OpenStudy (mckenzieandjesus):

Find a solution to the following system of equations. -3x+3y = -9 -3x + y = 7

OpenStudy (mckenzieandjesus):

(–5, –8) (5, 2) (–2, 2) (0, 7)

OpenStudy (mckenzieandjesus):

@texaschic101

OpenStudy (texaschic101):

the first course of action here is to eliminate a variable.....either x or y, it does not matter. I believe it would be easier to eliminate the x's. -3x + 3y = -9 -- multiply everything by -1 -3x + y = 7 ----------- 3x - 3y = 9 (result of multiplying by -1) -3x + y = 7 -----------add -2y = 16 -- divide by -2 y = -8 then sub -8 back in for y in either of the original equations to find x do you have any questions about what I just did ?

OpenStudy (texaschic101):

now if you wanted to eliminate the y's instead of the x's, you would have multiplied the second equation by -3.....then added the equations together. But either way, you should arrive at the same answer.

OpenStudy (mckenzieandjesus):

27?

OpenStudy (mckenzieandjesus):

or 24?

OpenStudy (mckenzieandjesus):

@texaschic101

OpenStudy (mckenzieandjesus):

@LegendarySadist can u help?

OpenStudy (anonymous):

Just sub y=8 back into -3x + y = 7 and solve for x.

OpenStudy (anonymous):

Whoops, that would be y=-8

OpenStudy (mckenzieandjesus):

so -11?

OpenStudy (mckenzieandjesus):

@LegendarySadist

OpenStudy (mckenzieandjesus):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hopefully you understand how y = -8 what you do from here is replace every copy of 'y' in an equation like -3x + y = 7 and solve for x -3x + y = 7 -3x + (-8) = 7 ... replace y with -8 -3x - 8 = 7 -3x - 8+8 = 7+8 ... Add 8 to both sides. -3x + 0 = 15 -3x = 15 -3x/(-3) = 15/(-3) ... Divide both sides by -3 to fully isolate x. 1x/1 = -5 1x = -5 x = -5

jimthompson5910 (jim_thompson5910):

So together, x = -5 and y = -8 forming the ordered pair (-5,-8) if you were to graph the two lines, the point (-5,-8) would be on both lines. It is the point of intersection between the two lines

OpenStudy (mckenzieandjesus):

Yes I understand and oh okay thank you

jimthompson5910 (jim_thompson5910):

ok great